# Quantum Speedups in Regret Analysis of Infinite Horizon Average-Reward Markov Decision Processes

### v2

@@ IV Algorithmic Framework
However, in the absence of knowledge pertaining to true $P$, we propose to solve an optimistic version of Eq. (12) that will update RL agent’s policy in epochs. Specifically, at the start of an epoch $e$, the following optimization problem will be solved:
@@
Observe that Eq. (16) increases the search space for policy in a carefully computed neighborhood of certain transition probability estimates $\hat{P}$ via Eq. (20) at the start of epoch $e$. Therefore, solving Eq. (16) is equivalent to obtaining an optimistic policy as it extends the search space of original problem Eq. (12). Furthermore, the computation of $\hat{P}_{e}$ is key to our framework design. Since performing quantum mean estimation to quantum samples would require measuring and results in quantum samples collapsing, each quantum sample can only be used once. Thus, for a specific state-action pair $(s_{i},a_{i})$ we construct an estimator $\hat{P}_{e}(\cdot|s_{i},a_{i})$ used in epoch $e$ that consists of a weighted average of the estimator $\hat{P}_{e-1}(\cdot|s_{i},a_{i})$ and the estimator obtained using the quantum samples in the latest completed epoch $\tilde{P}_{e}(\cdot|s_{i},a_{i})$ as follows:
@@
$\tilde{P}_{e}(s_{j}|s_{i},a_{i})=\texttt{SGE}(\{\bm{\psi}_{j}\}_{e}^{i},\nu_{e-1}(s_{i},a_{i}),\delta_{e})\quad\forall s_{j}\in\mathcal{S}$ (24)
$\{\bm{\psi}_{j}\}_{e}^{i}=\{\bm{\psi}^{t}_{j}\,|\;t_{e-1}^{start}\leq t\leq t_{e}^{start}-1,(s_{t},a_{t})=(s_{i},a_{i})\}$ (25)
where $t_{e}^{start}$ is the starting step of epoch $e$ and $\{\bm{\psi}_{j}\}_{e}^{i}$ being the quantum samples of the indicator $\mathbbm{1}[s^{\prime}=s_{j}|s_{i},a_{i}]$ obtained in the previous epoch with the total number being $\nu_{e-1}(s_{i},a_{i})$. Different from the works in classical approaches based on [35], our algorithm utilizes each quantum sample only once for estimating all $\hat{P}_{e}$ throughout the entire process. This strategy fits with our quantum framework since quantum samples would collapse once being measured. In Appendix A, we show that Eq. (16) is a convex optimization problem, therefore standard optimization solvers can be employed to solve it. Finally, the agent’s policy at the start of epoch $e$ is given by:
@@
Next, we formally present our method Q-UCRL in Algorithm IV.
{algorithm}
[ht] Q-UCRL
@@
11:     Observe $s_{t+1},r(s_{t},a_{t})$ and quantum samples $\bm{\psi}^{t}$.
@@
Algorithm IV proceeds in epochs, and each epoch $e$ contains multiple rounds of RL agent’s interaction with the environment. At each $t$, by playing action $a_{t}$ for insantaneous state $s_{t}$ through current epoch policy $\pi_{e}$, the agent collects next state configuration consisting of classical signals $s_{t+1},r(s_{t},a_{t})$ and a quantum sample $|\bm{\psi}\rangle^{(t)}$. Additionally, the agent keeps track of its visitations to the different states via the variables $\{\nu_{s,a},\mu_{s,a,s^{\prime}}\}$. Note that a new epoch is triggered at Line 14 of algorithm IV is aligned to the widely used concept of doubling trick used in model-based RL algorithm design. Finally, the empirical visitation counts and probabilities up to epoch $e$, i.e., $\{N_{e+1}(s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ attributes are updated (line 15-21), and the policy is updated via re-solving from Eq. (16) - (20) and using Eq. (26).
V Theoretical Results
In this section, we characterize the regret accumulated for $T$ rounds by running Q-UCRL in a unknown MDP hybrid Q-Environment. In the following, we first present a crucial result bounding the error accumulated till $e$ epochs between the true transition probabilities $\{P(\cdot|s,a)\}$ and the agent’s estimates $\{P_{e}(\cdot|s,a)\}$.
@@ V Theoretical Results
Execution of Q-UCRL (Algorithm IV) up to the beginning of epoch $e+1$ with total $e$ completed epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for $\tilde{P}_{e+1}$:
$\mathbb{P}[|\tilde{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq\frac{C\sigma\log(1/\delta)}{\nu_{e}(s,a)}]\geq 1-\delta$ (27)
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, where $C=c\log^{3/2}(T)\log\log(T)$ for some $c\in\mathbb{R}$ and $\{\nu_{e}(s,a)\}$ are the state-action pair visitation counts up in epoch $e$, as depicted in Line 12 of Algorithm IV, and $\{P(s^{\prime}|s,a),\tilde{P}_{e+1}(s^{\prime}|s,a)\}$ are the actual and estimated transition probabilities by Eq. (24).
@@
For all $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ , since $\tilde{P}_{e+1}$ is obtained by Eq. (24), following Lemma 1 would obtain:
$\mathbb{P}[|\tilde{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq\frac{\sigma\log(1/\delta)}{n}]\geq 1-\delta$ (28)
Where there exists $c\in\mathbb{R}$ such that
$\nu_{e}(s,a)\leq cn\log^{3/2}(T)\log\log(T)$ (29)
Thus, Eq. (28) can be transformed into Eq. (27) by substituting $n$ with $\nu_{e}{(s,a)}/C$
@@
Execution of Q-UCRL (Algorithm IV) up to the beginning of epoch $e+1$ with total $e$ completed epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for $\hat{P}_{e+1}$:
$\mathbb{P}[|\hat{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq\frac{eC\sigma\log(e/\delta)}{N_{e+1}(s,a)}]\geq 1-\delta$ (30)
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, where $C=c\log^{3/2}(T)\log\log(T)$ for some $c\in\mathbb{R}$ and $\{\nu_{e}(s,a)\}$ are the state-action pair visitation counts up in epoch $e$, as depicted in Line 12 of Algorithm IV, and $\{P(s^{\prime}|s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ are the actual and weighted average of transition probabilities by Eq. (21).
@@
Denote event $|\tilde{P}_{i}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq\frac{C\sigma\log(1/\delta)}{\nu_{i}(s,a)}$ as event $E_{i}$. Thus if events $E_{1},...,E_{e}$ all occur, we would have:
@@
$\leq\frac{\sum_{i=1}^{e}C\sigma\log(1/\delta)}{\sum_{j=1}^{e}\nu_{j}(s,a)},$ (35)
$=\frac{eC\sigma\log(1/\delta)}{\sum_{j=1}^{e}\nu_{j}(s,a)},$ (36)
$=\frac{eC\sigma\log(1/\delta)}{N_{e+1}(s,a)},$ (37)
@@
Execution of Q-UCRL (Algorithm IV) up to total $e$ epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for confidence parameter $\delta_{e}=\frac{1}{SAt_{e}^{7}}$ with probability at least $1-1/t_{e}^{6}$:
@@
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCeloge}{\max\{1,N_{e}(s,a)\}},$ (42)
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, where $\{N_{e-1}^{(s,a)}\}$ is the cumulative state-action pair visitation counts up to epoch $e$, as depicted in Line 17 of Algorithm IV, and $\{P(\cdot|s,a),P_{e}(\cdot|s,a)\}$ are the actual and estimated transition probabilities.
@@
$\mathbb{P}\left[|{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}|\leq\frac{Ce\log(e)}{N_{e}{(s,a)}}+Ce\epsilon\right]$
@@
$\mathbb{P}\left[|{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}|\geq\frac{Ce\log e+Ce\log(S^{2}At_{e}^{7})}{N_{e}(s,a)}\right]$
@@
Here, it is critical to highlight that we obtain this high probability bound for model errors by careful utilization of Lemma 1 and quantum signals $\bm{\psi}^{t}$ via Eq. (42) - (49), while avoiding classical concentration bounds for probability norms that use martingale style analysis such as in the work [36]. Finally, plugging the bound of term (a) as described by event $\mathcal{E}$, we obtain the following with probability $1-1/t_{e}^{6}$:
@@
$\leq\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}+\sum_{s^{\prime}\in\mathcal{S}}\frac{Ce\log e+Ce\log(S^{2}At_{e}^{7})}{\max\{1,N_{e}(s,a)\}},$ (50)
$\leq\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}+\sum_{s^{\prime}\in\mathcal{S}}\frac{eE\log e+7Ce\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}},$ (51)
$=\frac{7S(1+Ce)\log(S^{2}At_{e})+SCeloge}{\max\{1,N_{e}(s,a)\}},$ (52)
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCeloge}{\max\{1,N_{e}(s,a)\}}.$ (53)
@@
which essentially captures the gap in the average rewards by playing action $a$ in state $s$ accordingly to some policy $\pi$ in one step of the optimistic transition model $P_{e}$ w.r.t. the actual probabilities $P$. This leads us to the following key results that connect the regret expression in Eq. (61) and the Bellman errors.
@@
[ Lemma 5.2, [9]] The difference in long-term expected reward by playing optimistic policy of epoch $e$, i.e., $\pi_{e}$ on the optimistic transition model estimates, i.e., $\Gamma_{\pi_{e}}^{{P}_{e}}$, and the expected reward collected by playing $\pi_{e}$ on the actual model $P$, i.e., $\Gamma_{\pi_{e}}^{{P}}$ is the long-term average of the Bellman errors. Mathematically, we have:
@@
[Lemma 5.3, [9]] The Bellman errors for any arbitrary state-action pair $(s,a)$ corresponding to epoch $e$ is bounded as follows:
@@
where $\tilde{h}(\cdot)$ (defined in [32]) is the bias of the optimistic MDP $P_{e}$ for which the optimistic policy $\pi_{e}$ obtains the maximum expected reward in the confidence interval.
@@
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCeloge}{\max\{1,N_{e}(s,a)\}}\|\tilde{h}(\cdot)\|_{\infty},$ (65)
Consequently, it is necessary to bound the model bias error term in Eq. (65). We recall that the optimistic MDP $P_{e}$ corresponding to epoch $e$ maximizes the average reward in the confidence set as described in the formulation of $\tilde{\bm{\Xi}}_{e}$. In other words, $\Gamma_{\pi_{e}}^{P_{e}}\geq\Gamma_{\pi_{e}}^{P^{\prime}}$ for every $P^{\prime}$ in the confidence interval. Specifically, the model bias for optimistic MDP can be explicitly defined in terms of Bellman equation [32], as follows:
@@
[Lemma D.5, [9]] For an MDP with the transition model $P_{e}$ that generates rewards $r(s,a)$ on playing policy $\pi_{e}$, the difference of biases for states $s$ and $s^{\prime}$ are bounded as:
@@
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCeloge}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}.$ (68)
@@
In an unknown MDP environment $\mathcal{M}\triangleq(\mathcal{S},\mathcal{A},P,r,D)$, the regret incurred by Q-URL (Algorithm IV) across $T$ rounds is bounded as follows:
$\mathcal{R}_{[0,T-1]}=\mathcal{O}\Bigg{(}{S^{5}A^{4}T_{\texttt{mix}}\log^{11/2}(S^{2}AT)\log\log(T)}\Bigg{)}.$ (69)
@@
$\leq\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\frac{7S(1+Ce)\log(S^{2}AT)+SCeloge}{\max\{1,N_{e}(s_{t},a_{t})\}}T_{\texttt{mix}},$ (72)
@@
$\hskip 28.45274pt\frac{7S(1+Ce)\log(S^{2}AT)+SCeloge}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}\Big{)},$ (73)
@@
$\hskip 28.45274pt\frac{7S(1+Ce)\log(S^{2}AT)+SCeloge}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}},$ (74)
$=\sum_{s,a}T_{\texttt{mix}}(7S(E+CE^{2})\log(S^{2}AT)+SCE^{2}\log E)$
@@
$\leq\sum_{s,a}T_{\texttt{mix}}(7S(E+CE^{2})\log(S^{2}AT)+SCE^{2}\log E)E,$
$=T_{\texttt{mix}}E(7S^{2}A(E+CE^{2})\log(S^{2}AT)+S^{2}ACE^{2}\log E),$
$=\mathcal{O}\Bigg{(}{S^{5}A^{4}T_{\texttt{mix}}\log^{11/2}(S^{2}AT)\log\log(T)}\Bigg{)}.$ (76)
where in Eq. (71), we directly incorporate the result of Lemma 5. Next, we obtain Eq. (72) by using the Bellman error bound in Eq. (68), as well as unit sum property of occupancy measures [32]. Furthermore, in Eq. (72) note that we have omitted the regret contribution of bad event (Remark 1) which is atmost $\tilde{\mathcal{O}}\Bigg{(}\frac{2ST_{\texttt{mix}}}{T^{4}}\Bigg{)}$ as obtained below:
@@
Subsequently, Eq. (74) - (75) are obtaining by using the definition of epoch-wise state-action pair visitation frequencies $\nu_{e}(s,a)$ and the epoch termination trigger condition in Line 13 of Algorithm IV. Finally, we obtain Eq. (76) by the using Proposition 18 in Appendix C.2 of [8]. ∎
@@
We reiterate that our characterization of regret incurred by Q-UCRL in Theorem 1 is a martingale-free analysis and thus deviates from its classical counterparts [10, 8, 9].
VI Conclusions and Future Work

### v3

@@ 4 Algorithmic Framework
However, in the absence of knowledge pertaining to true $P$, we propose to solve an optimistic version of Eq. (11) that will update RL agent’s policy in epochs.
Remark 1.
Algorithm 2 runs in epochs wherein the agent’s action policy $\pi$ is updated only at the beginning of a new epoch, and the policy is held constant throughout the epoch.
Specifically, at the start of an epoch $e$, the following optimization problem will be solved:
@@
Observe that Eq. (15) increases the search space for policy in a carefully computed neighborhood of certain transition probability estimates $\hat{P}$ via Eq. (19) at the start of epoch $e$. Also in Eq. (19), recall that $S=|\mathcal{S}|$ is the number of states, $A=|\mathcal{A}|$ is the number of actions, $N_{e}(s,a)$ is the number of visitations of state-action pair $(s,a)$ upto the end of epoch $e$.
Note that, solving Eq. (15) is equivalent to obtaining an optimistic policy as it extends the search space of original problem Eq. (11). Furthermore, the computation of $\hat{P}_{e}$ is key to our framework design. Since performing quantum mean estimation to quantum samples would require measuring and results in quantum samples collapsing, each quantum sample can only be used once. We emphasize that the quantum mean estimator will only run at the end of the epoch, performing measurements to update the transition probability for the next epoch.
Details on Quantum Mean Estimator: At the end of the epoch, we use all $(s_{i},a_{i})$ observed in the samples, and perform quantum mean estimation on all the samples $|\bm{\psi}\rangle^{t}$ obtained for that $(s_{i},a_{i})$. More precisely, $\tilde{P}_{e}(\cdot|s_{i},a_{i})$ is the quantum mean estimation for the random variable $P_{e}(\cdot|s_{i},a_{i})$ performed by QBounded (Algorithm 1 in (?)) using the quantum samples $\{|\bm{\psi}\rangle\}^{e}_{i}$ observed in the previous epoch with the total number being $\nu_{e-1}(s_{i},a_{i})$:
${\{|\bm{\psi}\rangle\}^{e}_{i}=\{|\bm{\psi}\rangle^{t}\,|\;t_{e-1}^{start}\leq t\leq t_{e}^{start}-1,(s_{t},a_{t})=(s_{i},a_{i})\}}$ (20)
where $t_{e}^{start}$ is the starting step of epoch $e$. Using these samples, the transition probability is obtained as
${\tilde{P}_{e}(\cdot|s_{i},a_{i})=\texttt{QBounded}_{S}(\{P_{e}(\cdot|s_{i},a_{i}),L_{2},n_{e}^{s_{i},a_{i}},\delta_{e})}$ (21)
For a specific state-action pair $(s_{i},a_{i})$, we construct an estimator $\hat{P}_{e}(\cdot|s_{i},a_{i})$ used in epoch $e$ that consists of a weighted average of the estimator $\hat{P}_{e-1}(\cdot|s_{i},a_{i})$ and the estimator obtained using the quantum samples in the latest completed epoch $\tilde{P}_{e}(\cdot|s_{i},a_{i})$ as follows:
@@
In particular, given that $\nu_{e-1}(s_{i},a_{i})$ is the maximum number of quantum experiments we can perform in Eq. (21), we set
$n_{e}^{s_{i},a_{i}}=\left\lfloor\frac{\nu_{e-1}(s_{i},a_{i})}{c\log^{1/2}(T\sqrt{S})}\right\rfloor$ (25)
for some $c\in\mathbb{R}$ such that the term $cn\log^{1/2}(n\sqrt{S})$ is greater than the number of quantum experiments needed for performing Eq. (21), which is $\mathcal{O}\Big{(}n\log^{1/2}(n\sqrt{S})\Big{)}$. In addition, since $\mathbb{E}[||P_{e}(\cdot|s_{i},a_{i})||_{2}]\leq 1$, we set $L_{2}=1.$ Different from the works in classical approaches based on (?), our algorithm utilizes each quantum sample only once for estimating all $\hat{P}_{e}$ throughout the entire process. This strategy fits with our quantum framework since quantum samples would collapse once being measured. In Appendix A, we show that Eq. (15) is a convex optimization problem, therefore standard optimization solvers can be employed to solve it. Finally, the agent’s policy at the start of epoch $e$ is given by:
@@
Next, we formally present our method Q-UCRL in Algorithm 2.
Algorithm 2 Q-UCRL
@@
11:     Observe $s_{t+1},r(s_{t},a_{t})$ and quantum samples $|\bm{\psi}\rangle^{t}$.
@@
Algorithm 2 proceeds in epochs, and each epoch $e$ contains multiple rounds of RL agent’s interaction with the environment. At each $t$, by playing action $a_{t}$ for insantaneous state $s_{t}$ through current epoch policy $\pi_{e}$, the agent collects next state configuration consisting of classical signals $s_{t+1},r(s_{t},a_{t})$ and a quantum sample $|\bm{\psi}\rangle^{(t)}$. Additionally, the agent keeps track of its visitations to the different states via the variables $\{\nu_{s,a},\mu_{s,a,s^{\prime}}\}$. Note that a new epoch is triggered at the “if” block of Line 15 in algorithm 2, which in turn is aligned to the widely used concept of doubling trick used in model-based RL algorithm design. Finally, the empirical visitation counts and probabilities up to epoch $e$, i.e., $\{N_{e+1}(s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ attributes are updated (line 15-21), and the policy is updated via re-solving from Eq. (15) - (19) and using Eq. (26).
5 Theoretical Results
In this section, we characterize the cumulative regret for $T$ rounds by running Q-UCRL in a unknown MDP hybrid Q-Environment. In the following, we first present a crucial result bounding the error accumulated till $e$ epochs between the true transition probabilities $\{P(\cdot|s,a)\}$ and the agent’s estimates $\{P_{e}(\cdot|s,a)\}$.
@@ 5 Theoretical Results
Execution of Q-UCRL (Algorithm 2) up to the beginning of epoch $e+1$ with total $e$ completed epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for $\tilde{P}_{e+1}$:
$\mathbb{P}[|\tilde{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq{\frac{C\log(S/\delta)}{\nu_{e}(s,a)}}]\geq 1-\delta$ (27)
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, where $C={c\log^{1/2}(T\sqrt{S})}$ for some $c\in\mathbb{R}$ defined in Eq. (25) and $\{\nu_{e}(s,a)\}$ are the state-action pair visitation counts up in epoch $e$, as depicted in Line 12 of Algorithm 2, and $\{P(s^{\prime}|s,a),\tilde{P}_{e+1}(s^{\prime}|s,a)\}$ are the actual and estimated transition probabilities by Eq. (21).
@@
For all $(s,a)\in\mathcal{S}\times\mathcal{A}$ , since $\tilde{P}_{e+1}$ is obtained by Eq. (21), we obtain using Lemma 1:
$\mathbb{P}[{||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{\log(S/\delta)}{n_{e+1}^{s_{i},a_{i}}}}]\geq 1-\delta,$ (28)
Thus, by substituting $n_{e}^{s_{i},a_{i}}$ with $\nu_{e}{(s,a)}/C$, Eq. (28) would produce:
${\mathbb{P}[||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{C\log(S/\delta)}{\nu_{e}(s,a)}]\geq 1-\delta}$ (29)
Eq. (29) can be transformed into Eq. (27) by taking the entry-wise expression.
@@
Execution of Q-UCRL (Algorithm 2) up to the beginning of epoch $e+1$ with total $e$ completed epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for $\hat{P}_{e+1}$:
$\mathbb{P}[|\hat{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq\frac{eC\log(e{S}/\delta)}{N_{e+1}(s,a)}]\geq 1-\delta$ (30)
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, where $C={c\log^{1/2}(T\sqrt{S})}$ for some $c\in\mathbb{R}$ and $\{\nu_{e}(s,a)\}$ are the state-action pair visitation counts up in epoch $e$, as depicted in Line 12 of Algorithm 2, and $\{P(s^{\prime}|s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ are the actual and weighted average of transition probabilities by Eq. (22).
@@
Denote event $|\tilde{P}_{i}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq{\frac{C\log(S/\delta)}{\nu_{e}(s,a)}}$ as event $E_{i}$. Thus if events $E_{1},...,E_{e}$ all occur, we would have:
@@
$\leq\frac{\sum_{i=1}^{e}C\log({S}/\delta)}{\sum_{j=1}^{e}\nu_{j}(s,a)},$ (35)
$=\frac{eC\log({S}/\delta)}{\sum_{j=1}^{e}\nu_{j}(s,a)},$ (36)
$=\frac{eC\log({S}/\delta)}{N_{e+1}(s,a)},$ (37)
@@
Execution of Q-UCRL (Algorithm 2) up to total $e$ epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for confidence parameter $\delta_{e}=\frac{1}{SAt_{e}^{7}}$ with probability at least $1-1/t_{e}^{6}$:
@@
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}},$ (42)
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, where $\{N_{e-1}^{(s,a)}\}$ is the cumulative state-action pair visitation counts up to epoch $e$, as depicted in Line 17 of Algorithm 2, and $\{P(\cdot|s,a),P_{e}(\cdot|s,a)\}$ are the actual and estimated transition probabilities.
@@
$\mathbb{P}\left[|{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}|\leq\frac{Ce\log(e{S})}{N_{e}{(s,a)}}+Ce\epsilon\right]$
@@
$\mathbb{P}\left[|{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}|\geq\frac{Ce\log{(eS)}+Ce\log(S^{2}At_{e}^{7})}{N_{e}(s,a)}\right]$
@@
Here, it is critical to highlight that we obtain this high probability bound for model errors by careful utilization of Lemma 1 and quantum signals $\bm{\psi}^{t}$ via Eq. (42) - (49), while avoiding classical concentration bounds for probability norms that use martingale style analysis such as in the work (?). Finally, plugging the bound of term (a) as described by event $\mathcal{E}$, we obtain the following with probability $1-1/t_{e}^{6}$:
@@
$\leq\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}+\sum_{s^{\prime}\in\mathcal{S}}\frac{Ce\log{(eS)}+Ce\log(S^{2}At_{e}^{7})}{\max\{1,N_{e}(s,a)\}},$ (50)
$\leq\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}+\sum_{s^{\prime}\in\mathcal{S}}\frac{Ce\log{(eS)}+7Ce\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}},$ (51)
$=\frac{7S(1+Ce)\log(S^{2}At_{e})+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}},$ (52)
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}.$ (53)
@@
which essentially captures the gap in the average rewards by playing action $a$ in state $s$ accordingly to some policy $\pi$ in one step of the optimistic transition model $P_{e}$ w.r.t. the actual probabilities $P$. Also, we note here that the definition of $Q_{\pi}^{{P}_{e}}$ function is associated to the value function definition in eq (5).
This leads us to the following key results that connect the regret expression in Eq. (61) and the Bellman errors.
@@
[ Lemma 5.2, (?)] The difference in long-term expected reward by playing optimistic policy of epoch $e$, i.e., $\pi_{e}$ on the optimistic transition model estimates, i.e., $\Gamma_{\pi_{e}}^{{P}_{e}}$, and the expected reward collected by playing $\pi_{e}$ on the actual model $P$, i.e., $\Gamma_{\pi_{e}}^{{P}}$ is the long-term average of the Bellman errors. Mathematically, we have:
@@
[Lemma 5.3, (?)] The Bellman errors for any arbitrary state-action pair $(s,a)$ corresponding to epoch $e$ is bounded as follows:
@@
where $\tilde{h}(\cdot)$ (defined in (?)) is the bias of the optimistic MDP $P_{e}$ for which the optimistic policy $\pi_{e}$ obtains the maximum expected reward in the confidence interval.
@@
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}\|\tilde{h}(\cdot)\|_{\infty},$ (65)
Consequently, it is necessary to bound the model bias error term in Eq. (65). We recall that the optimistic MDP $P_{e}$ corresponding to epoch $e$ maximizes the average reward in the confidence set as described in the formulation of eq (15). In other words, $\Gamma_{\pi_{e}}^{P_{e}}\geq\Gamma_{\pi_{e}}^{P^{\prime}}$ for every $P^{\prime}$ in the confidence interval. Specifically, the model bias for optimistic MDP can be explicitly defined in terms of Bellman equation (?), as follows:
@@
[Lemma D.5, (?)] For an MDP with the transition model $P_{e}$ that generates rewards $r(s,a)$ on playing policy $\pi_{e}$, the difference of biases for states $s$ and $s^{\prime}$ are bounded as:
@@
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}.$ (68)
@@
In an unknown MDP environment $\mathcal{M}\triangleq(\mathcal{S},\mathcal{A},P,r,D)$, the regret incurred by Q-URL (Algorithm 2) across $T$ rounds is bounded as follows:
$\mathcal{R}_{[0,T-1]}=\mathcal{O}\Bigg{(}{S^{5}A^{4}T_{\texttt{mix}}{\log^{3}(\frac{T}{SA})\log^{1/2}(T\sqrt{S})\log(S^{2}AT)}}\Bigg{)}.$ (69)
@@
$\leq\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s_{t},a_{t})\}}T_{\texttt{mix}},$ (72)
@@
$\hskip 28.45274pt\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}\Big{)},$ (73)
@@
$\hskip 28.45274pt\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}},$ (74)
$=\sum_{s,a}T_{\texttt{mix}}(7S(E+CE^{2})\log(S^{2}AT)+SCE^{2}\log{(ES)})$
@@
$\leq\sum_{s,a}T_{\texttt{mix}}(7S(E+CE^{2})\log(S^{2}AT)+SCE^{2}\log{(ES)})E,$
$=T_{\texttt{mix}}E(7S^{2}A(E+CE^{2})\log(S^{2}AT)+S^{2}ACE^{2}\log{(ES)}),$
$=\mathcal{O}\Bigg{(}{S^{5}A^{4}T_{\texttt{mix}}{\log^{3}(\frac{T}{SA})\log^{1/2}(T\sqrt{S})\log(S^{2}AT)}}\Bigg{)}.$ (76)
where in Eq. (71), we directly incorporate the result of Lemma 5. Next, we obtain Eq. (72) by using the Bellman error bound in Eq. (68), as well as unit sum property of occupancy measures (?). Furthermore, in Eq. (72) note that we have omitted the regret contribution of bad event (Remark 2) which is atmost $\tilde{\mathcal{O}}\Bigg{(}\frac{2ST_{\texttt{mix}}}{T^{4}}\Bigg{)}$ as obtained below:
@@
Subsequently, Eq. (74) - (75) are obtaining by using the definition of epoch-wise state-action pair visitation frequencies $\nu_{e}(s,a)$ and the epoch termination trigger condition in Line 13 of Algorithm 2. Finally, we obtain Eq. (76) by the using Proposition 18 in Appendix C.2 of (?). ∎
@@
We reiterate that our characterization of regret incurred by Q-UCRL in Theorem 1 is a martingale-free analysis and thus deviates from its classical counterparts (?, ?, ?).
6 Conclusions and Future Work